A ball of mass 200g, traveling with velocity of 100mls collide with another ball of mass 800g moving at 59ml in the same direction. If they stick together, what will be their common velocity?


1
Expert's answer
2021-02-28T07:32:27-0500

We can find the common velocity of two balls from the law of conservation of momentum:


"m_1v_1+m_2v_2=(m_1+m_2)v_c,""v_c=\dfrac{m_1v_1+m_2v_2}{m_1+m_2},""v_c=\dfrac{0.2\ kg\cdot100\ \dfrac{m}{s}+0.8\ kg\cdot59\ \dfrac{m}{s}}{0.2\ kg+0.8\ kg}=67.2\ \dfrac{m}{s}."